By Pete Maddicks and Tony Howard
Tony H has again penned this month’s update, thank you Tony.
Monday 6th March.
Bob L, Charles L, Steve T, Tony H.
First working day of March. We found the cab roof in the paintshop. We were advised that the sides would need the paint scraping of to the bare metal. A start was made doing this job.
The vacuum pipe that goes from the brake in the cab, through the floor and connects to the train pipe was put in place to make sure it did not interfere with the drop grate mechanism. Luckily it did not. Then during lunchtime it started to rain, and as the loco is presently outside that put paid to any further work.
Monday 13th March.
Bob L, Charles L, Peter M, Steve T, Tony H.
Arrived at Bridgnorth today to find the loco still outside. As it raining quite heavily it did not look as though we could get any work done. Whithin 30 minutes the rain had stopped and it turned out to be quite pleasant.
Charles and Steve sorted out the the vacuum pipe that joins the vacuum retaining valve to the train pipe, it is now a good fit. It meant having to put an extra flange in. Tony found the extra long bolts to fix this all together.
The cab roof which is now in the paint shop was raised up onto four blocks of wood and Bob gave the sides a coat of paint. A length of 2" pipe was obtained for the next phase of vacuum pipe work.
Monday 20th March.
Bob L, Charles L, Dave Mc, Peter M, Phil H, Steve T, Tony H.
On arrival at Bridgnorth the loco was still outside. A word to the right person, and the loco was shunted into the shed and over the wheel drop pit. This was ideal for sorting out the vacuum pipe. The pipe that we acquired last week had to have the outside diameter reduced so the special nut that fixes the vacuum pipe to the brake valve will fit. A lot more fettling was needed on the new brass fittings. By late afternoon they more or less fitted correctly including the water gauge glass.
Saturday 25th March.
Kevin S, Peter M, Simon B, Tony H.
Today's job was the drop grate operation mechanism. We have the parts that we ordered a few weeks ago. These have been laser cut. We now looked for some off cut 60mm and 40mm round bar. We did find what we needed a little oversize so wiĺ need to go on the lathe to reduce the diameter. More photos and measurements were taken to make certain we were doing the job correctly.
Monday 27th March.
Bob L, Bryn T, Charles L, Dave Mc, Peter M, Phil H, Steve T, Tony H.
There was no lathe available so Phil took the material to do the turning at home. More work was done by Charles to the modifications of the vacuum pipe.
Tony and Steve finally finished the fettling of the water gauge frame and also made two templates for gaskets for the gauge frame and the blower valve. Bob and Bryn got more paint on the smokebox. Dave and Peter got the smokebox door into the paint shop.